# Probabilities of Chat LLMs Are Miscalibrated but Still Predict Correctness on Multiple-Choice Q&A

### v3

@@ Abstract
We study 15 large language models (LLMs) fine-tuned for chat and find that their maximum softmax probabilities (MSPs) are consistently miscalibrated on multiple-choice Q&A. However, those MSPs might still encode useful uncertainty information. Specifically, we hypothesized that wrong answers would be associated with smaller MSPs compared to correct answers. Via rigorous statistical testing, we show that this hypothesis holds for models which perform well on the underlying Q&A task. We also find a strong direction correlation between Q&A accuracy and MSP correctness prediction, while finding no correlation between Q&A accuracy and calibration error. This suggests that within the current fine-tuning paradigm, we can expect correctness prediction but not calibration to improve as LLM capabilities progress. To demonstrate the utility of correctness prediction, we show that when models have the option to abstain, performance can be improved by selectively abstaining based on the MSP of the initial model response, using only a small amount of labeled data to choose the MSP threshold.
[…]
[rm]bfb
Probabilities of Chat LLMs Are Miscalibrated but Still Predict Correctness on Multiple-Choice Q&A
Benjamin Plautthanks: Corresponding author UC Berkeley plaut@berkeley.edu                        Nguyen X. Khanh UC Berkeley kxnguyen@berkeley.edu                        Tu Trinh UC Berkeley tutrinh@berkeley.edu
@@ 1 Introduction
Large language models (LLMs) have demonstrated profound capabilities in many domains, but continue to sometimes generate plausible-sounding false responses (Huang et al., 2023). In one high-profile case, an LLM-based system invented a litany of nonexistent court cases, leading to formal sanctions for two lawyers (Mangan, 2023). Although ongoing work has reduced the rate of these mistakes, [^1] See, for example, https://huggingface.co/spaces/hallucinations-leaderboard/leaderboard. LLMs will inevitably face situations that surpass the boundaries of their existing knowledge. In those situations, it is unrealistic to expect these models (or any intelligent agents, including humans) to always make perfect decisions. Rather than confidently misleading users, LLMs should be able to detect unfamiliar situations and act cautiously (e.g., decline to answer).
In this paper, we study whether LLMs can determine the correctness of their own answers to multiple-choice questions. If so, this would directly enable LLMs to decline to answer when they are likely to be incorrect. Prior work has fine-tuned LLMs with labeled data to recognize questions beyond their knowledge (Kadavath et al., 2022; Zhang et al., 2023), but this approach is costly and could interfere with other capabilities of the model. More broadly, such approaches do not answer the fundamental scientific question of whether LLMs already possess the necessary information to decide when to abstain. That is, can this ability be invoked rather than learned?
We investigate the maximum softmax probability (MSP) as a potential source of innate uncertainty information. Specifically, our goal is to understand the relationship between the MSP of an LLM response and the correctness of that response.
We evaluate 15 LLMs fine-tuned for chat (henceforth “chat LLMs” for brevity) on five different Q&A datasets. Figure 1 shows a sample question prompt. The selected LLMs cover a range of sizes, capabilities and architectures, and include both open-source and proprietary models. To our knowledge, our work is the most comprehensive study on LLM correctness-awareness. This comprehensiveness bolsters the robustness of our claims but also enables cross-model comparisons which reveal novel insights.
@@
Calibration. We first ask whether the MSP is calibrated (DeGroot and Fienberg, 1983; Nguyen and O’Connor, 2015), meaning that among responses with an MSP of $p\%$, $p\%$ are correct. Calibrated MSPs enable fully unsupervised abstention policies with theoretical guarantees: a calibrated model that answers only when the MSP is higher than $1-\varepsilon$ guarantees that the chance of an incorrect answer is at most $\varepsilon$. However, we show that this approach is not generally viable for chat LLMs. Figure 2 shows that most LLMs in our study have poorly calibrated MSPs. In particular, the MSPs are consistently overconfident. Furthermore, improved performance on the underlying Q&A task does not necessarily translate to better calibration: Figure 3 shows no correlation between overall Q&A accuracy (percentage of questions answered correctly) and calibration error ($p=0.32$). This is one of the cross-model comparisons mentioned above.
@@
Through rigorous statistical testing, we demonstrate that the MSPs of chat LLMs can indeed predict correctness. [^2] Correction prediction is measured by the Area Under the Receiver Operating Characterstic curve (AUROC). See Section 5 for details. Moreover, the predictiveness is stronger for models which perform better on the underlying Q&A task ($p<10^{-3}$). This second cross-model comparison suggests that the ability to predict correctness will strengthen as the general capabilities of LLMs improve (e.g., by scaling up data and model sizes). The same is not true of calibration, as discussed above. These contrasting results reveal a novel dichotomy between two approaches to uncertainty quantification.
Q&A with abstention. In addition to demonstrating the predictive power of the MSP and maximum logit, we provide a proof-of-concept for how this information can be leveraged to reduce LLM harm in practice. We analyzed a variant of the original Q&A task where models can also abstain and receive 1 point per correct answer, 0 points per abstention, and $-c$ points per wrong answer. We found that for both $c=1$ and $c=2$, selectively abstaining based on the MSP and/or maximum logit led to substantial improvements over the base models. We used a mere 20 data points (i.e., 20 randomly selected questions and their answers) to select the abstention threshold.
In summary, we show the following:
@@
    A small amount of labeled data is enough to translate correctness prediction into an effective abstention method.
The paper proceeds as follows. Section 2 discusses related work and Section 3 describes our general experimental setup. Sections 4, 5, and 6 present our calibration results, AUROC results, and Q&A-with-abstention results, respectively.
@@ 2 Related work
The first is that they primarily studied raw pretrained (i.e., not fine-tuned) LLMs. In particular, their well-known finding that LLMs are well-calibrated applies only to raw pretrained models. In this way, our work complements theirs by showing that their finding of good calibration fails to generalize to LLMs fine-tuned for chat. (They actually briefly analyzed one fine-tuned model and found that it is quite miscalibrated.) Correctness-awareness may also be more important for fine-tuned models, since the casual user is less likely to interact with raw pretrained LLMs.
@@
The second is that they only studied MSP calibration and not MSP correctness prediction. [^3] They did study correctness prediction in the different context of training LLMs to abstain. We discuss those results separately below. This may be because good calibration directly yields theoretically grounded correctness prediction, so for raw pretrained models, good calibration may suffice. However, our finding that chat LLMs have miscalibrated MSPs motivates the separate question of whether MSPs can predict correctness.
@@
The third is comprehensiveness. They only tested a single series of models, while we test 6 series of models (or 8, depending on how one counts) and 15 models total. Our comprehensiveness crucially enables cross-model comparisons, as discussed in Section 1. In particular, we have statistical evidence that the correlation between correctness prediction and Q&A accuracy (and the lack of a correlation between calibration and Q&A accuracy) may extend to models that do not even exist yet. In contrast, it is harder to claim that the findings of Kadavath et al. (2022) generalize to other models, since they essentially have a sample size of one.
@@
Overall, our work complements theirs. Viewing our work and theirs side-by-side suggests that fine-tuning degrades calibration of LLMs and this effect is not mitigated as models become more capable. However, this procedure only distorts rather erases uncertainty information in LLMs, and that uncertainty information does become more useful as models become more capable.
LLM calibration. Several other papers have studied MSP calibration in LLMs (OpenAI, 2023; Krause et al., 2023; Tian et al., 2023; Zhou et al., 2024). These papers have found some evidence that chat LLMs produce overconfident MSPs, but the experimental conditions vary significantly between the papers and most of the papers study only a single model or family of models. To our knowledge, we are the first to present a comprehensive and unified evaluation of the MSP calibration of LLMs fine-tuned for chat.
Orthogonal to our work, Tian et al. (2023) studied the calibration of verbalized confidence estimates from LLMs. Although conceptually related, this is technically distinct from MSP calibration.
@@
Beyond LLMs. The MSP has been used for anomaly/out-of-distribution detection in a variety of other contexts, including coreference resolution tasks (Nguyen and O’Connor, 2015) pre-trained BERT models (Hendrycks et al., 2020), and image classification (Hendrycks et al., 2022; Hendrycks and Gimpel, 2017).
@@ 3 Experimental setup
This section presents our general experimental setup. Elements of the setup which are specific to calibration, correctness prediction, or abstention are discussed in Sections 4,  5 and 6, respectively. We have attached the code for all of our experiments and analysis as supplementary material.
@@
Multiple-choice Q&A. We chose to study multiple-choice Q&A because there is exactly one correct answer: this allows us to study the core hypothesis of whether MSPs are predictive of correctness without having to deal with more complex issues such as degrees of correctness or multiple valid phrasings of the same correct answer.
@@
Datasets. We based our experimental framework on the original Hugging Face LLM leaderboard (Beeching et al., 2023). We used all five multiple-choice Q&A datasets [^4] The leaderboard also includes the GSM8k dataset, which we excluded since it is not multiple-choice. from that leaderboard: ARC-Challenge (Clark et al., ), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022), and WinoGrande (Sakaguchi et al., 2021). We randomly sampled 6,000 questions from each dataset, except for those with fewer than 6,000, in which case we used all of the questions (ARC-Challenge and TruthfulQA have 2,590 and 817 questions, respectively). The 6,000 number was chosen to make the experiment duration manageable.
@@
Prompting style. To test our hypothesis in the simplest possible setting, we used a plain zero-shot prompting style. For comparison, we also ran the experiments with 5-shot prompting and obtained similar results (Appendix B).
We also used two different phrasings to ensure that our results were not an artifact of the specific phrasing. One phrasing is shown in Figure 1 and the other appears in Appendix A.
Models. We tested 15 LLMs: 13 open-source and two proprietary. The open-source LLMs were chosen based on a combination of performance on the aforementioned leaderboard and the number of downloads on Hugging Face. The open-source models we selected are Falcon (7B and 40B) (Almazrouei et al., 2023), Llama 2 (7B, 70B) (Touvron et al., 2023), Llama 3 (8B, 70B) and Llama 3.1 (8B, 70B) (AI@Meta, 2024), Mistral 7B v0.2 (Jiang et al., 2023), Mixtral 8x7B (Jiang et al., 2024), SOLAR 10.7B (Kim et al., 2023), and Yi (6B and 34B) (01-ai, 2023).
We used the fine-tuned “chat" or “instruct" versions of all models. Our focus on fine-tuned models is due to the unreliability of the non-fine-tuned base models. We were unable to get many of the base models to consistently respond in the correct format, despite significant prompting effort.
@@
All of the open-source LLMs were accessed through the Hugging Face interface and were run with dynamic 4-bit quantization, which has been shown to preserve performance while significantly reducing computational requirements (Dettmers et al., 2023). The experiments on open-source LLMs took lasted 1000 GPU-hours using NVIDIA RTX A6000 GPUs.
@@
We also tested two proprietary LLMs for which we could obtain softmax probabilities through an API: OpenAI’s GPT-3.5 Turbo (Brown et al., 2020; Ouyang et al., 2022) and GPT-4o (OpenAI, ; OpenAI, 2023). The OpenAI API does not provide pre-softmax logits, so we could not compute Max Logit AUROCs for those two models. [^5] We also could not test “reasoning” models such as o1 and o3-mini, since the API provides neither logits nor softmax probabilities for those models. These experiments took about a day and cost about \$120.
@@
3.1 Computing MSP and Max Logit
Let $V$ be a set of tokens (a vocabulary) and let $\boldsymbol{x}$ be a sequence of tokens from $V$. For each token $y\in V$ and prefix $\boldsymbol{x}$, an LLM computes a logit $L(y;\boldsymbol{x})$. It then applies a softmax function to the logits to derive the probability of $y$ being the next token:
$P(y\mid\boldsymbol{x})=\frac{\exp(L(y;\boldsymbol{x}))}{\sum_{z\in V}\exp(L(z;\boldsymbol{x}))}$
The Max Logit of token $y$ is $\max_{y\in V}L(y;\boldsymbol{x})$ and the MSP is $\max_{y\in V}P(y\mid\boldsymbol{x})$. Note that the Max Logit and MSP both correspond to the same token: $\operatorname*{arg\,max}_{y\in V}L(y;\boldsymbol{x})=\operatorname*{arg\,max}_{y\in V}P(y\mid\boldsymbol{x})$. In our experiments, we formulated each question as a prompt (Figure 1), then used greedy decoding to generate a response (i.e., we always picked the token with the maximum logit).
For typical supervised learning classification tasks such as those studied in Hendrycks and Gimpel (2017) and Hendrycks et al. (2022), there is a single MSP and maximum logit per response. However, an LLM response can consist of multiple tokens and the model produces an MSP and Max Logit for each token.
Ideally, we would like extract from the response a single token that indicates the LLM’s answer to the question. For this reason, our prompts state that the models should reply in the format “A.”/“B.”/etc. Fortunately, nearly all responses were in the correct format: 87% of responses started with “A.”/“B.”/etc, and 92% contained “A.”/“B.”/etc somewhere in the response. Those numbers increase to 93% and 97% respectively when the weaker models are excluded (Falcon, Llama 2, Mistral, Yi 6B).
As such, we simply searched the output string for the first occurrence of “A.”/“B.”/etc and then recorded the MSP and Max Logit corresponding to that capital letter token, since that capital letter token determines the LLM’s answer. If there was no occurence of “A.”/“B.”/etc , we searched for just “A”/“B”/etc, although as mentioned, this case was rare. The same search process was used for both computing MSP/Max Logit and evaluating correctness. If the search failed, we recorded the MSP and Max Logit as zero, and labeled the LLM’s response as wrong. [^7] We considered excluding these responses, but we reasoned that practically speaking, they are incorrect. This was also not an impactful decision, as only 0.4% of responses were unparseable. This number drops to 0.1% when the Falcon, Llama 2, Mistral, and Yi 6B models are excluded.
To verify that our results are not sensitive to the specifics of this search process, we performed a secondary analysis where we replaced the MSP of the answer token with the product of MSPs across all tokens. Those results were similar and can be found in Appendix C.

### v4

@@ Abstract
We evaluate 15 large language models (LLMs) fine-tuned for chat on multiple-choice Q&A. Consistent with prior work, we find that their maximum softmax probabilities (MSPs) are consistently miscalibrated on multiple-choice Q&A. However, those MSPs might still encode useful uncertainty information. Specifically, we hypothesized that wrong answers would be associated with smaller MSPs compared to correct answers. Via rigorous statistical testing, we show that this hypothesis holds for models which perform well on the underlying Q&A task. We also find a strong direct correlation between Q&A accuracy and MSP correctness prediction, while finding no correlation between Q&A accuracy and calibration error. This suggests that within the current fine-tuning paradigm, we can expect correctness prediction but not calibration to improve as LLM capabilities progress. To demonstrate the utility of correctness prediction, we show that when models have the option to abstain, performance can be improved by selectively abstaining based on the MSP of the initial model response, using only a small amount of labeled data to choose the MSP threshold.
@@ 1 Introduction
Large language models (LLMs) have demonstrated profound capabilities in many domains, but still sometimes generate plausible-sounding false responses (Huang et al., 2023). In one high-profile case, an LLM-based system invented a litany of nonexistent court cases, leading to formal sanctions for two lawyers (Mangan, 2023). Although ongoing work has reduced the rate of these mistakes, [^1] See, for example, https://huggingface.co/spaces/hallucinations-leaderboard/leaderboard. LLMs will inevitably face situations that surpass the boundaries of their existing knowledge. In those situations, it is unrealistic to expect these models (or any intelligent agents, including humans) to always make perfect decisions. Rather than confidently misleading users, LLMs should be able to detect unfamiliar situations and act cautiously (e.g., decline to answer).
In this paper, we study whether LLMs can determine the correctness of their own answers to multiple-choice questions. If so, this would directly enable LLMs to decline to answer when they are likely to be incorrect. One natural uncertainty metric for neural networks is the maximum softmax probability (MSP). We study a simple but fundamental question: what does the MSP of an LLM response tell us about whether that response is correct? We are far from the first to study this question, but prior findings are dispersed over a variety of contexts and experimental setups. Our goal is to provide a comprehensive, unified study that both reproduces prior findings and enables novel insights.
1.1 Basic setup
We evaluate 15 LLMs fine-tuned for chat (henceforth “chat LLMs” for brevity) on five different Q&A datasets. Figure˜1 shows a sample question prompt. The selected LLMs cover a range of sizes, capabilities, and architectures, and include both open-weight and proprietary models. To our knowledge, our work is the most comprehensive study of LLM correctness-awareness.
In multiple-choice Q&A, the MSP is defined as follows. We first compute the probability that the LLM assigns to each answer token (e.g., “A”, “B”, “C”, etc.) and renormalize those probabilities to sum to 1. The MSP is then the maximum of those probabilities. The LLM’s response to the question is the token corresponding to the MSP. This approach is consistent with Biderman et al. (2024); Chen et al. (2025); Zhao et al. (2021), among others.
@@ 1.1 Basic setup
Calibration. We first ask whether the MSP is calibrated (DeGroot & Fienberg, 1983; Nguyen & O’Connor, 2015), meaning that among responses with an MSP of $\%, pp\%$ are correct. Calibrated MSPs enable fully unsupervised abstention policies with theoretical guarantees: a calibrated model that answers only when the MSP is higher than $1-\varepsilon$ guarantees that the chance of an incorrect answer is at most $\varepsilon$. However, prior work has shown that this approach is not generally viable for chat LLMs (OpenAI, 2023; Zhu et al., 2023): in particular, the MSPs are consistently overconfident.
Our investigation begins by validating those findings on a more diverse set of models and tasks. Furthermore, the comprehensiveness of our study enables robust cross-model comparisons that were not possible in prior evaluations that only considered a few models. In particular, we find that calibration does not improve as chat models become more capable.
@@
Through rigorous statistical testing, we demonstrate that the MSPs of chat LLMs can indeed predict correctness. [^2] Correctness prediction is measured by the Area Under the Receiver Operating Characteristic curve (AUROC). This finding is not surprising given that calibration can often be restored with appropriate rescaling of MSPs (see Section˜2 for further discussion). Once again, the more interesting finding is the cross-model comparison. We find that this predictiveness is stronger for models which perform better on the underlying Q&A task ($p<10^{-3}$). In other words, the ability to predict correctness will likely strengthen as the general capabilities of LLMs improve (e.g., by scaling up data and model sizes). The same is not true of calibration, as discussed above. These contrasting results reveal a fundamental dichotomy between two approaches to uncertainty quantification, summarized by Table˜1.
Q&A with abstention. In addition to demonstrating the predictive power of the MSP and maximum logit, we provide a proof-of-concept for how this information can be leveraged to reduce LLM harm in practice. We analyzed a variant of the original Q&A task where models can also abstain and receive 1 point per correct answer, 0 points per abstention, and c-c points per wrong answer. We found that for both $c=1$ and $c=2$, selectively abstaining based on the MSP and/or maximum logit led to substantial improvements compared to never abstaining. We used a mere 20 data points per dataset (i.e., 20 randomly selected questions and their answers) to select the abstention threshold.
Base models. Although our focus is chat LLMs, we run the same experiments for base (i.e., non-fine-tuned) models. Consistent with prior work (Kadavath et al., 2022), we find that the base models are much better calibrated than the chat models, and the calibration of base models does improve as Q&A accuracy improves. These positive calibration results suggest that the MSP will also predict correctness, which we confirm rigorously.
Table 1: A summary of the results in our paper. Capability is measured by Q&A accuracy, calibration is measured by expected calibration error, and correctness prediction is measured by AUROC. Some of these findings are more novel than others; see Section˜2 for details.
Chat LLMs Base LLMs
Calibrated? ✗ (Figure˜2, left) ✓ (Figure˜6, left)
Calibration improves with capability?
✗  (Figure˜2, right) ✓ (Figure˜6, right)
MSP predicts correctness? ✓ (Table˜2) ✓ (Table˜8)
MSP correctness prediction improves with capability?
✓ (Figure˜3) ✓ (Figure˜7)
In summary, our key results are the following:
@@
    A small amount of labeled data can translate correctness prediction into an effective abstention method.
The paper proceeds as follows. Section˜2 discusses related work and Section 3 covers our general experimental setup. Sections 4, 5, and 6 present our results for chat LLMs on calibration, correctness prediction, and Q&A-with-abstention, respectively. Section˜7 covers the analogous results for base models. Appendix˜A provides more detail on the above results. The remaining appendices cover 5-shot prompting experiments (which produce similar results), other measures of uncertainty (margin and entropy), a post-hoc calibration experiment, and dataset-level analysis.
@@ 2 Related work
The first is that they primarily study base LLMs. In particular, their well-known findings that (1) LLMs are well-calibrated and (2) calibration improves further as model size/capability increases apply only to base models. In this way, our work complements theirs by showing that their finding of good calibration fails to generalize to LLMs fine-tuned for chat. Correctness-awareness may also be more important for fine-tuned models, since the casual user is less likely to interact with base LLMs.
@@
The second is that they only studied MSP calibration and not MSP correctness prediction. [^3] They did study correctness prediction in the different context of training LLMs to abstain. We discuss those results separately below. This may be because good calibration directly yields theoretically grounded correctness prediction, so for base models, good calibration may suffice. However, our finding that chat LLMs have miscalibrated MSPs motivates the separate question of whether MSPs can predict correctness.
@@
The third is comprehensiveness. They only tested a single series of models, while we test 6 series of models (or 8, depending on how one counts) and 15 models total. Our comprehensiveness crucially enables cross-model comparisons, as discussed in Section˜1. In particular, we have statistical evidence that the correlation between correctness prediction and Q&A accuracy (and the lack of a correlation between calibration and Q&A accuracy) may extend to models that do not even exist yet. In contrast, it is harder to claim that the findings of Kadavath et al. (2022) generalize to other models, since they essentially have a sample size of one.
@@
Overall, our work complements theirs. Viewing our work and theirs side-by-side suggests that fine-tuning degrades calibration of LLMs and this effect is not mitigated as models become more capable. However, this procedure only distorts rather than erases uncertainty information in LLMs, and that uncertainty information does become more useful as models become more capable.
LLM calibration. Uncertainty quantification in LLMs is a very active area and a full survey is beyond the scope of this paper; we direct the interested reader to Geng et al. (2024). Several prior papers have found evidence that chat LLMs are miscalibrated (He et al., 2023; OpenAI, 2023; Zhu et al., 2023). These papers also showed that the calibration curves of chat LLMs are roughly monotone (e.g., Figure˜2 left), which suggests – but does not prove – that the MSP may predict correctness even though it is miscalibrated.
However, these papers generally only study one model (or at most a few models), and each paper uses different experimental conditions. In contrast, we present a comprehensive and unified evaluation of the MSP calibration and correctness-awareness of 15 chat LLMs. We also mention the simultaneous work by Xiao et al. (2025), which studies the calibration of four chat LLMs and finds the same overconfidence.
Post-hoc calibration. The miscalibration of chat LLMs has motivated the development of methods to restore calibration by rescaling the MSP in some way. Temperature scaling and variants thereof are especially popular, which have been shown to improve calibration without harming performance (Shen et al., 2024; Xie et al., 2024). There also exist generation-based methods. For example, Zhang et al. (2024) replace the model’s chosen response with “All other options are wrong” and test whether the model still selects that option. Ulmer et al. (2024) trained a predictor based on input-output pairs and calibration targets. Xiao et al. (2025) proposed a calibration-aware fine-tuning method applied after standard fine-tuning which restores calibration.
The success of post-hoc calibration methods aligns with our finding that MSPs can predict correctness even when poorly calibrated. In both cases, the key insight is that although the fine-tuning process disrupts calibration, the MSPs of chat LLMs retain an underlying uncertainty signal which can be recovered.
Verbalized uncertainty in LLMs. An alternative way to obtain uncertainty estimates from LLMs is to prompt them directly. One benefit of this approach is that it requires no access to the internals of the model. However, this approach has produced mixed results: LLMs can sometimes verbalize calibrated confidence levels (Lin et al., 2022a; Tian et al., 2023), but can also be highly overconfident (Xiong et al., 2024). Interestingly, Xiong et al. (2024) found that LLMs typically state confidence values in the range of 80-100%, usually in multiples of 5, potentially in imitation of how humans discuss confidence levels. Nevertheless, prompting strategies remain an important tool for uncertainty quantification, along with measures based on the internal state (such as MSP).
@@
Beyond LLMs. The MSP has been used for anomaly/out-of-distribution detection in a variety of other contexts, including coreference resolution tasks (Nguyen & O’Connor, 2015), pre-trained BERT models (Hendrycks et al., 2020), and image classification (Hendrycks et al., 2022; Hendrycks & Gimpel, 2017).
@@ 3 Experimental setup
This section presents our general experimental setup. Elements of the setup that are specific to calibration, correctness prediction, or abstention are discussed in Sections 4, 5, and 6, respectively. Our code can be found at https://github.com/bplaut/llm-calibration-and-correctness-prediction.
@@
Multiple-choice Q&A. We chose to study multiple-choice Q&A because there is exactly one correct answer. This allows us to study the core hypothesis of whether MSPs are predictive of correctness without potential confounders such as degrees of correctness or multiple valid phrasings of the same correct answer.
@@
Datasets. We based our experimental framework on the original Hugging Face LLM leaderboard (Beeching et al., 2023). We used all five multiple-choice Q&A datasets [^4] The leaderboard also includes the GSM8k dataset, which we excluded since it is not multiple-choice. from that leaderboard: ARC-Challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022b), and WinoGrande (Sakaguchi et al., 2021). We randomly sampled 6,000 questions from HellaSwag, MMLU, and WinoGrande. ARC-Challenge and TruthfulQA only have 2,590 and 817 questions, respectively, so we used all of those questions. The 6,000 number was chosen to make the experiment duration manageable.
@@
Prompting style. To test our hypothesis in the simplest possible setting, we used a plain zero-shot prompting style. For comparison, we also ran the experiments with 5-shot prompting and obtained similar results (Appendix B). We also used two different phrasings to ensure that our results were not an artifact of the specific phrasing. One phrasing is shown in Figure˜1 and the other appears in Appendix A.
@@
Models. We tested 15 LLMs: 13 open-weight and two proprietary. The open-weight LLMs were chosen based on a combination of performance on the aforementioned leaderboard and the number of downloads on Hugging Face. The open-weight models we selected are Falcon (7B and 40B) (Almazrouei et al., 2023), Llama 2 (7B, 70B) (Touvron et al., 2023), Llama 3.0 (8B, 70B) and Llama 3.1 (8B, 70B) (AI@Meta, 2024), Mistral 7B v0.2 (Jiang et al., 2023), Mixtral 8x7B (Jiang et al., 2024), SOLAR 10.7B (Kim et al., 2023), and Yi (6B and 34B) (01-ai, 2023). All of the open-weight LLMs were accessed through the Hugging Face interface and were run with dynamic 4-bit quantization, which has been shown to preserve performance while reducing computational requirements (Dettmers et al., 2023). We tested both fine-tuned “chat” and non-fine-tuned “base” versions of these 13 open-weight models. Unless we specify otherwise, the reader should assume we are referring to the “chat” versions of these models. The experiments on open-weight LLMs took about 2000 GPU-hours using NVIDIA RTX A6000 GPUs.
@@
We also tested two proprietary LLMs for which we could obtain softmax probabilities through an API: OpenAI’s GPT-3.5 Turbo (Brown et al., 2020; Ouyang et al., 2022) and GPT-4o (OpenAI, 2024). The OpenAI API does not provide pre-softmax logits, so we could not compute Max Logit AUROCs for those two models. [^5] We also could not test “reasoning” models such as o1 and o3, since the API provides neither logits nor probabilities for those models. We also did not have access to non-fine-tuned counterparts of these models. These experiments took about a day and cost about \$110.
@@
Computing MSP and Max Logit. Let $\mathcal{V}$ be a set of tokens (a vocabulary) and let $\boldsymbol{x}$ be a sequence of tokens from $\mathcal{V}$. For each token $y\in\mathcal{V}$ and prefix $\boldsymbol{x}$, an LLM computes a logit $L(y\mid\boldsymbol{x})$. A softmax function is applied to the logits to derive the probability of $y$ being the next token:
$P(y\mid\boldsymbol{x})=\frac{\exp(L(y\mid\boldsymbol{x}))}{\sum_{z\in\mathcal{V}}\exp(L(z\mid\boldsymbol{x}))}$
In our experiments, we formulated each question as a prompt $\boldsymbol{x}$ as in Figure 1. Let $\mathcal{T}$ be the set of possible answer tokens, i.e., $\mathcal{T}=\{A,B,\cdots\}$. We then computed $P(y\mid\boldsymbol{x})$ and $L(y\mid\boldsymbol{x})$ for each $y\in\mathcal{T}$, i.e., the probability and logit for $y$ being the first token of the response. [^7] A few models always begin responses with a dummy token, so for those models, we used the second token of the response. The LLM’s answer is the token with the maximum such probability and logit, i.e., $\operatorname*{arg\,max}_{y\in\mathcal{T}}P(y\mid\boldsymbol{x})=\operatorname*{arg\,max}_{y\in\mathcal{T}}L(y\mid\boldsymbol{x})$. The MSP and Max Logit are the probability and logit associated with that token, respectively. The probabilities were also renormalized to sum to 1 over tokens in $\mathcal{T}$. Formally,
$\text{MSP}(\boldsymbol{x})=\frac{\max_{y\in\mathcal{T}}P(y\mid\boldsymbol{x})}{\sum_{y\in\mathcal{T}}P(y\mid\boldsymbol{x})}\quad\quad\quad\quad\text{Max Logit}(\boldsymbol{x})=\max_{y\in\mathcal{T}}L(y\mid\boldsymbol{x})$